# Mitigating Structural Noise in Low-Resource S2TT: An Optimized Cascaded Nepali-English Pipeline with Punctuation Restoration

### v1

@@ 4 Results and Findings
4.1 Component-Level Performance
4.1.1 ASR Performance
The fine-tuned Wav2Vec2 XLS-R model was evaluated on the OpenSLR-54 test set (see Table 7).
Table 7: ASR Performance on OpenSLR-54 Test Set
@@ 4.1.1 ASR Performance
The resulting WER of 16.82% and a particularly low CER of 2.72% confirm the viability of the self-supervised transformer approach for low-resource Nepali ASR. This CER represents a substantial improvement over previous established supervised models on the same dataset (e.g., Dhakal et al. (2022) at 17.07% CER and Banjara et al. (2020) at 27.72% CER).
4.1.2 NMT Benchmarking
@@
Table 8: Benchmarking of the Final MarianMT Model Against Standard Test Sets and Previous Best Results
Test Set Metric Our Score Previous Best Score
@@
Tatoeba BLEU 39.66 3.5 (Base MarianMT mul-en)
chrF++ 55.73 0.168 (Base MarianMT mul-en)
@@
The NMT model achieved significant gains across all benchmarks, surpassing the previously reported state-of-the-art results Duwal and Bal (2019) on FLORES-200 by $\sim 8$99 BLEU points and drastically outperforming the base multilingual MarianMT model. This performance ensures that the NMT component is highly proficient when provided with high-quality, fully punctuated source text.
4.2 Quantifying Punctuation Impact on NMT
@@
Table 9: Impact of Punctuation on Translation Quality Across Test Sets (MarianMT Model Evaluation)
@@
Unpunctuated (Simulated ASR) 28.40 51.16 -11.26 -4.57
Figure 2: Comparative performance of NMT on punctuated and unpunctuated (Simulated ASR) inputs across three test sets. A consistent degradation is observed, confirming that punctuation loss substantially impacts translation quality.
4.3 End-to-End(S2TT) Evaluation
@@
4.3.1 Automatic Metrics (BLEU and chrF++)
@@
Figure 3: End-to-End S2TT Performance Comparison (BLEU & chrF++)
Table 10: Impact of System Configuration on Translation Quality
@@
4.3.2 Human Evaluation (Adequacy and Fluency)
@@
Table 11: Impact of System Configuration on Translation Fluency and Adequacy
Scenario Avg. Fluency (1–5) Avg. Adequacy (1–5)
A (Baseline) 3.677 3.581
B (Proposed - Full Pipeline) 3.774 3.628
C (Optimal Pipeline) 3.804 3.673
@@
  • Scenario B: ASR $\rightarrow$ Punctuation Restoration with space removal $\rightarrow$ NMT.
Figure 4: Human evaluation of S2TT output. Average fluency and adequacy scores (1–5 scale) for Scenarios A, B, and C.
Scenario C consistently yields the highest-quality output, achieving the best scores for both Fluency (3.804) and Adequacy (3.673). This result validates that the BLEU/chrF++ gains translate directly into a perceptibly higher-quality, more readable, and more meaning-preserving translations. Three independent human evaluators, and their consistent judgements reinforce the reliability of these findings.
4.3.3 Performance Analysis by Sentence Type
@@
Table 12: Sentence-Type-wise Human Evaluation Results
@@
Statements Avg. Adequacy 3.943 3.839 3.962
Statements Avg. Fluency 3.967 3.921 4.048
Commands Avg. Adequacy 3.167 3.421 3.416
Commands Avg. Fluency 2.981 3.171 3.226
Questions Avg. Adequacy 3.365 3.368 3.538
Questions Avg. Fluency 3.599 3.649 3.781
Named Entities Avg. Adequacy 3.269 3.887 3.269
Named Entities Avg. Fluency 3.259 3.921 3.323
Complex Avg. Adequacy 3.181 3.257 3.438
Complex Avg. Fluency 3.416 3.479 3.611
@@ 5 Discussion
5.1 Validation of Component-Level Refinements
@@
5.2 Quantifying Error Propagation
@@
5.3 Analysis of End-to-End Pipeline Performance
@@
Scenario B (ASR $\rightarrow$ Punctuation Restoration w/ removing spaces from input $\rightarrow$ NMT) performed sub-optimally, achieving lower score than Scenario C. This outcome demonstrates a crucial trade-off: the deliberate pre-processinng step of removing all spaces (token segmentation) and forcing the PRM to re-segment the entire utterance introduced new, cascading segmentation errors. These self-inflicted errors proved more detrimental to the downstream NMT model than the benefit gained from the restored punctuation alone. This finding suggests that for Nepali, preserving the ASR’s inherent word segmentation is critical and should not be discarded in favor of concurrent text-to-text re-segmentation unless the PRM is highly trained on a large-scale high-quality and diverse dataset for that specific task.
5.4 Validation by Human Assessment
The results from the human evaluation strongly corroborate the automatic metrics. Scenario C consistently scored the highest in both Fluency (3.804) and Adequacy (3.673). This correlation validates that the 4.90 BLEU gain is not merely a statistical artifact but translates directly into a perceivably higher-quality, more readable, and meaning-preserving translation for the end-user.
@@ 5.4 Validation by Human Assessment
5.5 Primary Contribution and Future Work
5.5.1 Primary Contribution
@@ 6 Conclusion
To mitigate this core problem, we proposed and validated a Punctuation Restoration Module (PRM) as a crucial intermediate component. The optimized pipeline configuration (Scenario C), which deployed the PRM on the raw ASR output to restore structural cues, yielded a significant performance increase, achieving a 4.90 BLEU point gain over the direct ASR-to-NMT baseline. This result was further validated by human evaluation, with the optimized system scoring highest in both Adequacy (3.673) and Fluency (3.804).

### v2

@@ 4. Results and Findings
4.1. Component-Level Performance
4.1.1. ASR Performance
The fine-tuned Wav2Vec2 XLS-R model was evaluated on the OpenSLR-54 test set (see Table 7). The resulting WER of 16.82% and a particularly low CER of 2.72% confirm the viability of the self-supervised transformer approach for low-resource Nepali ASR. This CER represents a substantial improvement over previous established supervised models on the same dataset (e.g., (Dhakal et al., 2022) at 17.07% CER and (Banjara et al., 2020) at 27.72% CER).
Table 7. ASR Performance on OpenSLR-54 Test Set
@@ 4.1.1. ASR Performance
4.1.2. NMT Benchmarking
@@
Table 8. Benchmarking of the Final MarianMT Model Against Standard Test Sets and Previous Best Results
Test Set Metric Our Score Previous Best
@@
Tatoeba BLEU 39.66 3.5a
chrF++ 55.73 0.168a
@@
The NMT model achieved significant gains across all benchmarks, surpassing the previously reported state-of-the-art results (Duwal and Bal, 2019) on FLORES-200 by $\sim 8$99 BLEU points and drastically outperforming the base multilingual MarianMT model. This performance ensures that the NMT component is highly proficient when provided with high-quality, fully punctuated source text.
4.2. Quantifying Punctuation Impact on NMT
@@
Table 9. Impact of Punctuation on Translation Quality Across Test Sets (MarianMT Model Evaluation)
@@
Unpunctuated (ASR) 28.40 51.16 -11.26 -4.57
Figure 2. Comparative performance of NMT on punctuated and unpunctuated (Simulated ASR) inputs across three test sets. A consistent degradation is observed, confirming that punctuation loss substantially impacts translation quality.
4.3. End-to-End(S2TT) Evaluation
@@
4.3.1. Automatic Metrics (BLEU and chrF++)
@@
Figure 3. End-to-End S2TT Performance Comparison (BLEU & chrF++)
Table 10. Impact of System Configuration on Translation Quality
@@
4.3.2. Human Evaluation (Adequacy and Fluency)
@@
Table 11. Impact of System Configuration on Translation Fluency and Adequacy with Inter-Rater Reliability ($\alpha$)
Scenario Avg. Fluency $\alpha$ (Fluency) Avg. Adequacy $\alpha$ (Adequacy)
A (Baseline) 3.677 0.723 3.581 0.783
B (Proposed) 3.774 0.746 3.628 0.780
C (Optimal) 3.804 0.748 3.673 0.785
@@
Scenario B: ASR $\rightarrow$ Punctuation Restoration (with space removal) $\rightarrow$ NMT.
Scenario C: ASR $\rightarrow$ Punctuation Restoration (without space removal) $\rightarrow$ NMT.
Figure 4. Human evaluation of S2TT output. Average fluency and adequacy scores (1–5 scale) for Scenarios A, B, and C.
Scenario C consistently yields the highest-quality output, achieving the best scores for both Fluency (3.804) and Adequacy (3.673). This result validates that the BLEU/chrF++ gains translate directly into a perceptibly higher-quality, more readable, and more meaning-preserving translations. Inter-rater reliability was measured using Krippendorff’s Alpha($\alpha$) for both metrics, yielding high agreement ($\alpha\geq 0.723$) across all systems, reinforcing the reliability of these findings.
4.3.3. Performance Analysis by Sentence Type
@@
Table 12. Sentence-Type-wise Human Evaluation Results
@@
Statements Adequacy 3.943 3.839 3.962
Statements Fluency 3.967 3.921 4.048
Commands Adequacy 3.167 3.421 3.416
Commands Fluency 2.981 3.171 3.226
Questions Adequacy 3.365 3.368 3.538
Questions Fluency 3.599 3.649 3.781
Named Entities Adequacy 3.269 3.887 3.269
Named Entities Fluency 3.259 3.921 3.323
Complex Adequacy 3.181 3.257 3.438
Complex Fluency 3.416 3.479 3.611
@@ 5. Discussion
5.1. Validation of Component-Level Refinements
@@
5.2. Quantifying Error Propagation
@@
5.3. Analysis of End-to-End Pipeline Performance
@@
Scenario B (ASR $\rightarrow$ Punctuation Restoration with removing spaces from input $\rightarrow$ NMT) performed sub-optimally, achieving lower score than Scenario C. This outcome demonstrates a crucial trade-off: the deliberate pre-processinng step of removing all spaces (token segmentation) and forcing the PRM to re-segment the entire utterance introduced new, cascading segmentation errors. These self-inflicted errors proved more detrimental to the downstream NMT model than the benefit gained from the restored punctuation alone. This finding suggests that for Nepali, preserving the ASR’s inherent word segmentation is critical and should not be discarded in favor of concurrent text-to-text re-segmentation unless the PRM is highly trained on a large-scale high-quality and diverse dataset for that specific task.
5.4. Validation by Human Assessment
The results from the human evaluation strongly corroborate the automatic metrics. Scenario C consistently scored the highest in both Fluency (3.804) and Adequacy (3.673). This correlation validates that the 4.90 BLEU gain is not merely a statistical artifact but translates directly into a perceivably higher-quality, more readable, and meaning-preserving translation for the end-user. This is further supported by a high degree of consensus among evaluators, with a Krippendorff’s Alpha ($\alpha$) of 0.748 for fluency and 0.785 for adequacy across the optimal configuration.
@@ 5.4. Validation by Human Assessment
5.5. Primary Contribution and Future Work
5.5.1. Primary Contribution
@@ 6. Conclusion
To mitigate this core problem, we proposed and validated a Punctuation Restoration Module (PRM) as a crucial intermediate component. The optimized pipeline configuration (Scenario C), which deployed the PRM on the raw ASR output to restore structural cues, yielded a significant performance increase, achieving a 4.90 BLEU point gain over the direct ASR-to-NMT baseline. This result was further validated by human evaluation, with the optimized system scoring highest in both Adequacy (3.673) and Fluency (3.804), reinforced by inter-rater reliability ($\alpha\geq 0.72$) for all scenarios.